I am two numbers,the sum of my two numbers is 100, the difference in my 2 numbers is 24 what am i
Answer by opeodu(24)

Let the 1st and 2nd number be x and y. x+y=100 x-y=24 adding both equation gives 2x=124 x=62 substitute x in the 1st equation x+y=100 62+y=100 y=38